OpenStudy (anonymous):

The current in a solenoid with 23 turns per centimeter is 0.55 A . The solenoid has a radius of 1.6 cm . A long, straight wire runs along the axis of the solenoid, carrying a current of 14 A Find the magnitude of the net magnetic field a radial distance of 0.80 cm from the straight wire. Please I need help with this as soon as possible So far I calculated the magnetic field of the solenoid and the straight wire but the answer was wrong !:(

OpenStudy (irishboy123):

reckon it should look like this : \(B_{net} = \mu_o\sqrt{(n' I_s)^2 + (\dfrac{I_w}{2 \pi r})^2 }\) giving \( (1.26)10^{ -6} \sqrt{[(23)(0.55)]^2 + (\dfrac{14}{2 \pi (0.008)})^2 }\) recognise any of that ?!?!